OpenStudy (anonymous):

Mark deposited $1,200 in a new bank savings account at the beginning of the quarter. The principal earns 5% interest compounded quarterly. He made no other transactions during the period. How much should be in his account at the end of 6 months? Show your work. Remember: (3/12) = 0.25

OpenStudy (ranga):

5% interest compounded quarterly. The interest rate per quarter is 5/4% = 1.25% = 0.0125 Account balance at the end of 6 months = ? 6 months is two quarters. So we need to compound the interest twice. After the end of one quarter the amount will be: 1200 * (1 + 0.0125) = $1215 At the end of two quarters, the amount will be: 1215 * (1.0125) = $1230.19

OpenStudy (anonymous):

Thank you so much, this make so much more sense now :)

OpenStudy (ranga):

Alright! You are welcome.

OpenStudy (anonymous):

wow, ranga, i think you are really smart. youre a smart

OpenStudy (ranga):

You are too kind halai. :)